2000 tickets were sold at the ASU Thanksgiving Homecoming game. Student tickets cost $10 and non-student tickets cost $20.00. A total of $35,000 was collected. How many of each type of ticket were sold?
Answer by TimothyLamb(4379)

x = student tickets
y = non-student tickets
---
x + y = 2000
10x + 20y = 35000
---
put the system of linear equations into standard form
---
x + y = 2000
10x + 20y = 35000
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = student tickets = 500
y = non-student tickets = 1500
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---